Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $3,000. It seems4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo4. Assume that it is customary in the industry to bid jobs at 150% of total manufacturing cost (direct materials, direct labo
Answer #1
1
a. Plant wide overhead rate=Estimated Manufacturing overhead/Estimated direct labor=865200/618000=1.4=140% of direct labor cost
b. Amount of manufacturing overhead applied to Koopers job=Direct labor cost*140%=11400*140%=$ 15960
2
a. Departmental overhead rate=Estimated Manufacturing overhead of the department/Estimated direct labor cost of the department
Fabricating Machining Assembly
Estimated manufacturing overhead (1) 360500 412000 92700
Estimated direct labor cost 2 206000 103000 309000
Departmental overhead rate (1)/(2) 1.75 4.00 0.30
175% 400% 30%
b. Amount of manufacturing overhead applied to Koopers job=Direct labor cost of department*Departmental overhead rate
Cutting Machining Assembly
Direct labor cost 1 4000 600 6800
Departmental overhead rate 2 175% 400% 30%
Manufacturing overhead applied (1*2) 7000 2400 2040
4
a. Bid price:
$
Total direct material cost 5900
Total direct labor cost 11400
Manufacturing overhead applied (Refer 1-b.) 15960
Total manufacturing cost 33260
Bid price=150%*Total manufacturing cost=33260*150%=$ 49890
b. Bid price:
Cutting Machining Assembly Total
Direct material cost 3600 300 2000 5900
Direct labor cost 4000 600 6800 11400
Manufacturing overhead applied (Refer 2-b.) 7000 2400 2040 11440
Total manufacturing cost 28740
Bid price=150%*Total manufacturing cost=28740*150%=$ 43110